# A Comprehensive Survey on Multi-modal Conversational Emotion Recognition with Deep Learning

### v1

@@ Abstract.
Multi-modal conversation emotion recognition (MCER) aims to recognize and track the speaker’s emotional state using text, speech, and visual information in the conversation scene. Analyzing and studying MCER issues is significant to affective computing, intelligent recommendations, and human-computer interaction fields. Unlike the traditional single-utterance multi-modal emotion recognition or single-modal conversation emotion recognition, MCER is a more challenging problem that needs to deal with more complex emotional interaction relationships. The critical issue is learning consistency and complementary semantics for multi-modal feature fusion based on emotional interaction relationships. To solve this problem, people have conducted extensive research on MCER based on deep learning technology, but there is still a lack of systematic review of the modeling methods. Therefore, a timely and comprehensive overview of MCER’s recent advances in deep learning is of great significance to academia and industry. In this survey, we provide a comprehensive overview of MCER modeling methods and roughly divide MCER methods into four categories, i.e., context-free modeling, sequential context modeling, speaker-differentiated modeling, and speaker-relationship modeling. In addition, we further discuss MCER’s publicly available popular datasets, multi-modal feature extraction methods, application areas, existing challenges, and future development directions. We hope that our review can help MCER researchers understand the current research status in emotion recognition, provide some inspiration, and develop more efficient models.
@@ 1. Introduction
With the development of the mobile Internet, social media has become the main platform for people to communicate with each other (Park et al., 2016). Users can fully express their emotions through multi-modal data such as text, voice, image, and video. Building an multi-modal conversational emotion recognition model through multi-modal data is of vital practical significance for understanding users’ true emotional intentions (Ghosh and Anwar, 2021). Therefore, researchers have been trying to give machines the ability to understand emotions in recent years (Li et al., 2022a), (Zhu et al., 2021).
Before the rise of multi-modal conversation emotion recognition research, early conversational emotion recognition methods (Kim, 2014; Tzinis et al., 2018; Zhong et al., 2019; Lotfian and Busso, 2019) mainly focused on single-modal text or speech, and they only considered the context dependencies between texts (or speech) and the semantic information of words (or audio) themselves to analyze and recognize emotions (Meng et al., 2021), (Ying et al., 2021), (Shou et al., 2022b). However, only extracting the emotional information contained in the text data may not be enough for the model to understand the emotion expressed by the speaker, because the speaker may be more reserved when expressing his opinion (Zhu et al., 2023; Ghosal et al., 2019). For example, a speaker may be veiled in expressing his anger, which may result in a more neutral utterance. In response to the above problems, multi-modal conversational emotion recognition (MCER) technology was proposed to solve the problem of insufficient expression of text semantic information (Shou et al., 2023b), (Shou et al., 2023a), (Ying et al., 2021). As shown in Fig. 1, MCER aims to extract semantic information complementary within and between modalities and identify the emotions expressed by speakers in text, audio, and video. The advantage of MCER is that when the emotional polarity of text information is insufficiently expressed (Lian et al., 2023), the model can use visual information (such as facial expressions, etc.) and audio information (such as tone, etc.) to enhance the emotional understanding of text information (Yin et al., 2022), (Yin et al., 2023a), (Yin et al., $5$), (Yin et al., 2023b).
@@
However, unlike traditional single-utterance multi-modal emotion recognition or single-modal conversation emotion recognition, MCER is a more challenging issue that requires consideration of factors such as multi-modal context, dialogue scenarios, speaker’s own emotional inertia, and interlocutor stimulation (Chudasama et al., 2022; Zhang et al., 2020). Powerful deep learning technology (Shou et al., 2022a) enables MCER to recognize emotion by fusing semantic features with complex emotional interactions. Feature fusion in MCER mainly considers intra-modal contextual semantic information fusion and inter-modal complementary semantic information fusion (Huang et al., 2020). On the one hand, intra-modal contextual semantic information fusion refers to extracting the temporal and spatial dependencies of speaker feature representations in each modality. On the other hand, complementary semantic information fusion between modalities refers to using the interactive information between different modalities to enhance the emotional understanding ability of the model. MCER synergistically improves the effect of emotion recognition by fusing the characteristics of various modal data, which has important theoretical significance for processing and understanding multi-modal data (Yang et al., 2022; Hou et al., 2023; Zhang et al., 2023a).
@@
Despite more and more people being devoted to researching new models and methods for multi-modal conversation emotion recognition (Ghosal et al., 2018; Chudasama et al., 2022; Liu et al., 2022), there is still a lack of understanding of the theoretical and methodological classification of multi-modal conversational emotion recognition, especially multi-modal conversational emotion recognition based on deep learning. To the best of our knowledge, this survey is the first comprehensive survey focusing on deep learning in multi-modal conversation emotion recognition. Existing surveys mainly focus on single-utterance multi-modal emotion recognition (Zhu et al., 2023) or single-modal conversation emotion recognition (Deng and Ren, 2021). Therefore, with the rapid penetration of deep learning in various fields, multi-modal conversation emotion recognition methods based on deep learning have become a research hotspot and require timely and comprehensive investigation.
Based on the above analysis, our survey summarizes the research work on multi-modal conversational emotion recognition. We first list some publicly available and popular datasets in the field of MCER, and list the commonly used feature extraction methods for each modality. Secondly, as shown in Fig. 2, we roughly divide the MCER methods into four categories, i.e., context-free modeling, sequential context modeling, distinguishing-speaker modeling, speaker-relationship modeling. Third, we provide evaluation indicators for MCER experiments. Fourth, we give the applications and problems of MCER. Finally, we illustrate future research directions.
@@
  • Abundant Resources: We collect relevant resour-ces about MCER, including state-of-the-art models and publicly available datasets. This paper can serve as a practical guide for learning and developing different emotion recognition algorithms.
@@ 2.1. IEMOCAP
The IEMOCAP dataset (Busso et al., 2008) was released in 2008 and contains 12.46 hours of conversations. The IEMOCAP dataset contains three modal features, i.e., video, audio and text, and it is the first multi-modal dataset for MCER. In the IEMOCAP dataset, ten theater actors express specific emotion categories (i.e., sad, neutral, frustrated, anger, happy, excited) through binary dialogue. To ensure the consistency and accuracy of annotation, each sentence is annotated by multiple experts.
@@ 2.2. MELD
The MELD dataset (Poria et al., 2019) is from the classic TV series Friends, which contains text, video and audio data. The MELD dataset contains a total of 13,709 video clips, and each sentence is labeled as a specific emotion (i.e., anger, neutral, fear, disgust, surprise, joy, disgust). In addition, the MELD dataset is also annotated by neutral, negative and positive three-category emotion. To ensure the consistency and accuracy of annotation, each sentence is annotated by multiple experts.
@@ 2.5. SEMAINE
SEMAINE (McKeown et al., 2011) is a multi-modal conversation data set, which contains four binary conversations between robots and humans. The SEMAINE data set has 95 dialogues with a total of 5798 sentences. Four emotional dimensions are marked: Valence, Arousal, Expectancy, and Power. Valence, Arousa, and Expectancy are continuous values in the range [-1, 1], and the size of the SEMAINE data set is small.
@@ 2.7. EmoContext
The EmoContext data set (Chatterjee et al., 2019) only contains text data. It has a total of 38,421 dialogues and a total of 115,263 sentences. Three types of emotions are marked: happiness, sadness, and anger. Although the EmoContext data is large, it is rarely used in conversational emotion recognition because it only contains text data.
@@ 3.1.1. Brief History of Conversational Emotion Recognition
The emotion recognition method based on the dictionary is the earliest used for emotion recognition (Hardeniya and Borikar, 2016), which motivated early work on naive bayes method (Frank et al., 2000). With the widespread application of machine learning algorithms in classification tasks, the representative machine learning classification algorithm (SVM (Rozgić et al., 2012), (Hu et al., 2007), and binary decision tree (Cichosz and Slot, 2007), (Lee et al., 2011), (Liu et al., 2018b), etc) has also begun to shine in the field of emotion recognition. The above-mentioned method determines the category of emotion by learning the polarity and occurrence frequency of emotional words in the text, which is difficult to extract the semantic information and context information.
Encouraged by the success of CNNs in the field of computer vision (CV), CNNs began to be migrated to text classification tasks and received extensive research attention (Khare and Bajaj, 2020), (Kwon et al., 2021), (Kollias and Zafeiriou, 2020). In 2017, Poria et al. (Poria et al., 2017) used LSTM for the first time to resolve dependencies between contexts. Since then, improvements, extensions and applications of LSTMs and GRUs have increased (Hazarika et al., 2018b), (Majumder et al., 2019), (Hazarika et al., 2018a), (Rajamani et al., 2021). Until recently, many GNN-based methods (e.g., (Ghosal et al., 2019), (Ishiwatari et al., 2020), (Shen et al., 2021), (Li et al., 2023), (Zadeh et al., 2018b)) emerged. Apart from CNNs, RNNs, and GNNs, many alternative Transformer-based methods (e.g., (Zhong et al., 2019), (Li et al., 2020), (Zhu et al., 2021)) have been developed in the past decades. We detail the categories to which these algorithms belong in Section 4.
@@ 3.1.2. Multi-modal Conversational Emotion Recognition Versus Traditional Machine Learning
MCER methods based on traditional machine learning (Rozgić et al., 2012), (Cichosz and Slot, 2007), (Lee et al., 2011), (Liu et al., 2018b), (Lin and Wei, 2005), (Hu et al., 2007) are closely related to hand-extracted features, which have attracted increasing attention from the data mining and emotion recognition communities. These methods aim to learn the feature embeddings of raw data for subsequent downstream tasks such as classification and clustering. The classic conversational emotion recognition method based on machine learning is to use support vector machine to map emotional features to a hyperplane and classify them (Rozgić et al., 2012), (Bhavan et al., 2019). However, these methods require a large amount of high-quality labeled data.
@@ 3.1.3. Conversational Emotion Recognition Versus Convolutional Neural Network
The CNN-based emotion recognition methods (Lu et al., 2023), (Kwon et al., 2021), (Kollias and Zafeiriou, 2020) are the first deep learning method to solve the emotion classification problem historically (Kim, 2014). These CNN-based methods employ convolutional filters to extract semantic features of text so that the model can use supervised learning to understand the meaning of text. Similar to machine learning algorithms, CNN can also map emotional features into vector space through mapping functions. The difference is that this mapping function is learned in an end-to-end manner. Since the convolution kernel extracts local receptive field information, CNN cannot contain contextual semantic information.
@@ 3.1.4. Multi-modal Conversational Emotion Recognition Versus Recurrent Neural Network
The RNN-based emotion recognition methods (Ma et al., 2019), (Tao and Liu, 2018), (Majumder et al., 2019), (Kollias and Zafeiriou, 2020) are developed on the basis of CNN, but they believe that contexts should be mutually influential and interdependent (Ma et al., 2019), (Tao and Liu, 2018). These RNN-based methods usually use LSTM or GRU (to avoid gradient disappearance or gradient explosion) to extract semantic features including context. Similar to CNN, RNN can also map emotional features into vector space through mapping functions in an end-to-end manner.
@@ 3.1.5. Multi-modal Conversational Emotion Recognition Versus Transformer
Similar to the RNN-based emotion recognition method, the Transformer-based emotion recognition methods (Huang et al., 2020), (Lian et al., 2021), (Tsai et al., 2019), (Rahman et al., 2020) also extract semantic information including context, and completes subsequent emotion classification based on this (Lian et al., 2021). However, unlike RNN, Transformer’s sequential context modeling ability is better than RNN. Therefore, the accuracy of the Transformer-based emotion recognition methods are significantly better than RNNs.

### v2

@@ Abstract.
Multi-modal conversation emotion recognition (MCER) aims to recognize and track the speaker’s emotional state using text, speech, and visual information. Compared with traditional single-utterance multi-modal emotion recognition or single-modal conversation emotion recognition, MCER is more challenging. It requires modeling complex emotional interactions and learning consistent and complementary semantics across multiple modalities. Although many deep learning-based approaches have been proposed for MCER, there is still a lack of systematic reviews summarizing existing modeling methods. Therefore, a timely and comprehensive overview of MCER’s recent advances in deep learning is of great significance. In this survey, we provide a comprehensive overview of MCER modeling methods and roughly divide MCER methods into four categories, i.e., context-free modeling, sequential context modeling, speaker-differentiated modeling, and speaker-relationship modeling. Unlike conventional taxonomies based on modality combinations or task-stage decomposition, our framework focuses on how models structurally capture conversational dynamics, speaker roles, and emotional dependencies. In addition, we further discuss MCER’s publicly available popular datasets, multi-modal feature extraction methods, application areas, existing challenges, and future development directions. We hope this review provides valuable insights into the current state of MCER research and inspires the development of more effective models.
@@ 1. Introduction
With the development of the mobile Internet, social media has become the main platform for people to communicate with each other (Park et al., 2016). Users can fully express their emotions through multi-modal data such as text, voice, image, and video. Building a multi-modal conversational emotion recognition model using multi-modal data is of vital practical significance for understanding users’ true emotional intentions (Ghosh and Anwar, 2021). Therefore, researchers have been trying to give machines the ability to understand emotions in recent years (Li et al., 2022a; Zhu et al., 2021; Zhang et al., 2023a).
@@
Before the emergence of multi-modal conversational emotion recognition, early methods (Kim, 2014; Tzinis et al., 2018; Zhong et al., 2019; Lotfian and Busso, 2019; Wen et al., 2024) primarily relied on single-modal data, such as text or speech. These approaches mainly focused on modeling contextual dependencies within the same modality and leveraging the semantic content of words or audio signals to recognize emotions (Meng et al., 2021; Ying et al., 2021; Shou et al., 2022; Ye et al., 2024b). However, relying solely on textual information may be insufficient for accurately interpreting a speaker’s emotional state, as speakers often express their opinions in a reserved or implicit manner (Zhu et al., 2023; Ghosal et al., 2019; Qin et al., 2024). For example, a speaker may be veiled in expressing his anger, which may result in a more neutral utterance. In response to the above problems, multi-modal conversational emotion recognition (MCER) technology was proposed to solve the problem of insufficient expression of text semantic information (Shou et al., 2023b, a; Ying et al., 2021; Qin et al., 2024). As shown in Fig. 1, MCER aims to extract semantic information complementary within and between modalities and identify the emotions expressed by speakers in text, audio, and video. One major advantage of MCER is its ability to enhance emotion understanding when the emotional polarity conveyed by text alone is insufficient (Lian et al., 2023). In such cases, the model can leverage visual cues (e.g., facial expressions) and acoustic signals (e.g., tone of voice) to supplement and enrich the emotional representation (Yin et al., 2022, 2023a, 5555; Yin et al., 2023b; Wang et al., 2024a). As illustrated in Fig. 2(a), the text modality alone predicts the speaker’s emotion to be “Neutral”, whereas the audio and visual modalities correctly predicts the speaker’s emotion to be “Sad”, highlighting the limitations of relying solely on text in emotional context inference. Furthermore, to validate the effectiveness of multimodal fusion, we provide LR-GCN latent space visualizations for both unimodal and multimodal settings in Fig. 2(b) and (c). It is evident that the multimodal feature space yields better inter-class separation, especially among subtle emotions such as “Neutral”, “Frustrated”, and “Sad”, demonstrating superior discriminative capability.
(a) Example of unimodal vs. multimodal in IEMOCAP.
(b) LR-GCN (unimodal)
(c) LR-GCN (multimodal)
Figure 2. Illustration of the advantage of multimodal fusion in emotion recognition. (a) Example from the IEMOCAP dataset showing that textual modality alone may fail to capture emotional intent (“Neutral”), while audio and visual modalities correctly identify the emotion as “Sad”. (b) Latent space visualization of GS-MCC with unimodal (text-only) input shows overlapping clusters and poor separation between emotion classes. (c) The same visualization under multimodal fusion shows significantly improved class separability, demonstrating the effectiveness of incorporating audio-visual information.
@@
However, unlike traditional single-utterance multimodal emotion recognition or single-modal conversation emotion recognition, MCER is a more challenging issue. It requires consideration of factors such as multimodal context, dialogue scenarios, the speaker’s emotional inertia, and the interlocutor’s stimulation (Chudasama et al., 2022; Zhang et al., 2020). Powerful deep learning technology (Zhang et al., 2024d) enables MCER to recognize emotion by fusing semantic features with complex emotional interactions. Feature fusion in MCER mainly considers intra-modal contextual semantic information fusion and inter-modal complementary semantic information fusion (Huang et al., 2020). On the one hand, intra-modal contextual semantic information fusion refers to extracting the temporal and spatial dependencies of speaker feature representations in each modality. On the other hand, complementary semantic information fusion between modalities refers to using the interactive information between different modalities to enhance the emotional understanding ability of the model. MCER synergistically improves the effect of emotion recognition by fusing the characteristics of various modal data, which has important theoretical significance for processing and understanding multi-modal data (Yang et al., 2022; Hou et al., 2023; Zhang et al., 2023a).
@@
Despite the growing number of researchers focusing on new models and methods for multimodal conversation emotion recognition (Ghosal et al., 2018; Chudasama et al., 2022; Liu et al., 2022), there is still a lack of understanding regarding the theoretical and methodological classification of multimodal conversational emotion recognition, particularly those based on deep learning. To the best of our knowledge, this survey is the first comprehensive survey focusing on deep learning in multi-modal conversation emotion recognition. Existing reviews mainly focus on multimodal emotion recognition with modal combination (Zhang et al., 2024e) or multimodal emotion recognition with task stage decomposition (i.e., feature extraction, feature fusion and classification) (Zhu et al., 2023), without fully considering conversational dynamics, speaker roles and emotional dependencies.
Different from previous taxonomies, we propose a novel classification framework that emphasizes how methods characterize and model conversational dynamics. Specifically, we categorize MCER methods into four distinct types: context-free modeling, sequential context modeling, speaker-differentiated modeling, and speaker-relationship modeling, as illustrated in Fig. 3.
Based on the above framework, this survey systematically reviews the key research efforts in MCER. First, we introduce several widely-used and publicly available datasets, along with commonly adopted feature extraction methods across modalities. Next, we detail the proposed modeling taxonomy and comprehensively analyze representative methods within each category. We then discuss evaluation metrics frequently used in MCER experiments. Following that, we examine real-world applications and key challenges faced in this field. Finally, we outline promising directions for future research.
@@
  • Abundant Resources: We collect relevant resources about MCER, including state-of-the-art models and publicly available datasets. This paper can serve as a practical guide for learning and developing different emotion recognition algorithms.
@@ 2.1. IEMOCAP
The interactive emotional dyadic motion capture database (IEMOCAP) dataset (Busso et al., 2008) was released in 2008 and contains 12.46 hours of conversations. The IEMOCAP dataset contains three modal features, i.e., video, audio and text, and it is the first multi-modal dataset for MCER. In the IEMOCAP dataset, ten theater actors express specific emotion categories (i.e., sad, neutral, frustrated, anger, happy, excited) through binary dialogue. To ensure the consistency and accuracy of annotation, each sentence is annotated by multiple experts.
@@ 2.2. MELD
The multimodal emotionLines (MELD) dataset (Poria et al., 2019) is from the classic TV series Friends, which contains text, video and audio data. The MELD dataset contains a total of 13,709 video clips, and each sentence is labeled as a specific emotion (i.e., anger, neutral, fear, disgust, surprise, joy, disgust). In addition, the MELD dataset is also annotated by neutral, negative and positive three-category emotion. To ensure the consistency and accuracy of annotation, each sentence is annotated by multiple experts.
@@ 2.5. SEMAINE
The sustained emotionally colored machine-human interaction (SEMAINE) (McKeown et al., 2011) is a multi-modal conversation data set, which contains four binary conversations between robots and humans. The SEMAINE data set has 95 dialogues with a total of 5798 sentences. Four emotional dimensions are marked: Valence, Arousal, Expectancy, and Power. Valence, Arousa, and Expectancy are continuous values in the range [-1, 1], and the size of the SEMAINE data set is small.
@@ 2.7. EmoContext
The emotion contextual detection (EmoContext) dataset (Chatterjee et al., 2019) only contains text data. It has a total of 38,421 dialogues and a total of 115,263 sentences. Three types of emotions are marked: happiness, sadness, and anger. Although the EmoContext data is large, it is rarely used in conversational emotion recognition because it only contains text data.
2.8. CH-SIMS
The Chinese single-label multimodal sentiment analysis (CH-SIMS) (Yu et al., 2020) is a benchmark dataset designed for Chinese multimodal sentiment analysis tasks. This dataset is constructed from real Chinese video conversations, covering three modalities: text, audio, and vision, and has a good foundation for multimodal fusion and alignment. CH-SIMS contains a total of about 10,000 Chinese sentence-level samples, all of which are accompanied by manually annotated continuous sentiment intensity labels (ranging from -1 to +1) and single-label classifications (positive, neutral, and negative). Compared with mainstream English multimodal sentiment analysis datasets such as IEMOCAP and MELD, CH-SIMS is more in line with the characteristics of Chinese language expression, especially when faced with semantic ambiguity (e.g., irony, sarcasm, etc.), there may be significant inconsistencies between modalities.
2.9. MuSE
Multimodal Sentiment dataset (MuSE) (Stappen et al., 2021) is a multilingual and multimodal emotion recognition dataset, which aims to study the performance of multimodal emotion modeling in different languages and cross-cultural contexts. The dataset contains natural speech videos from English and Spanish, with about 2,800 samples, covering three modalities: text, audio, and vision. Each video is recorded by real participants expressing freely around a specific topic, and provides continuous emotion labels, including valence, arousal, and dominance. The labels are derived from the fusion of self-reports and third-party manual evaluations to enhance the objectivity and consistency of annotations. At the modality level, MuSE provides high-quality speech features, expressions, and body posture information, and is equipped with precisely aligned text transcription data, making it suitable for research tasks such as multimodal fusion, modal alignment analysis, and modal inconsistency modeling. In addition, since the dataset has both multilingual and multimodal characteristics, it provides an important experimental platform for cross-lingual emotion transfer learning, multimodal collaborative modeling, and emotion recognition in low-resource languages.
@@ 3.1.1. Brief History of Conversational Emotion Recognition
The emotion recognition method based on the dictionary is the earliest used for emotion recognition (Hardeniya and Borikar, 2016), which motivated early work on naive bayes method (Frank et al., 2000). With the widespread application of machine learning algorithms in classification tasks, representative algorithms such as the support vector machine (SVM) (Rozgić et al., 2012; Hu et al., 2007) and binary decision tree (Cichosz and Slot, 2007; Lee et al., 2011; Liu et al., 2018b) have also gained prominence in emotion recognition. The above-mentioned method determines the category of emotion by learning the polarity and occurrence frequency of emotional words in the text, which is difficult to extract the semantic information and context information.
Encouraged by the success of convolutional neural networks (CNNs) in computer vision (CV), CNNs began to be migrated to text classification tasks and received extensive research attention (Khare and Bajaj, 2020; Kwon et al., 2021; Kollias and Zafeiriou, 2020). In 2017, Poria et al. (Poria et al., 2017) used long short-term memory (LSTM) for the first time to resolve dependencies between contexts. Since then, improvements, extensions and applications of LSTMs and gated recurrent units (GRUs) have increased (Hazarika et al., 2018b; Majumder et al., 2019; Hazarika et al., 2018a; Rajamani et al., 2021). Until recently, many graph neural networks (GNN)-based methods (e.g., (Ghosal et al., 2019; Ishiwatari et al., 2020; Shen et al., 2021; Li et al., 2023; Zadeh et al., 2018b)) emerged. Apart from CNNs, recurrent neural networks (RNNs), and GNNs, many alternative Transformer-based methods (e.g., (Zhong et al., 2019; Li et al., 2020; Zhu et al., 2021)) have been developed in the past decades. We detail the categories to which these algorithms belong in Section 4.
@@ 3.1.2. Multi-modal Conversational Emotion Recognition Versus Traditional Machine Learning
MCER methods based on traditional machine learning (Rozgić et al., 2012; Cichosz and Slot, 2007; Lee et al., 2011; Liu et al., 2018b; Lin and Wei, 2005; Hu et al., 2007) are closely related to hand-extracted features, which have attracted increasing attention from the data mining and emotion recognition communities. These methods aim to learn the feature embeddings of raw data for subsequent downstream tasks such as classification and clustering. The classic conversational emotion recognition method based on machine learning is to use support vector machine to map emotional features to a hyperplane and classify them (Rozgić et al., 2012; Bhavan et al., 2019). However, these methods require a large amount of high-quality labeled data.
@@ 3.1.3. Conversational Emotion Recognition Versus Convolutional Neural Network
The CNN-based emotion recognition methods (Lu et al., 2023; Kwon et al., 2021; Kollias and Zafeiriou, 2020) are the first deep learning method to solve the emotion classification problem historically (Kim, 2014). These CNN-based methods employ convolutional filters to extract semantic features of text so that the model can use supervised learning to understand the meaning of text. Similar to machine learning algorithms, CNN can also map emotional features into vector space through mapping functions. The difference is that this mapping function is learned in an end-to-end manner. Since the convolution kernel extracts local receptive field information, CNN cannot contain contextual semantic information.
@@ 3.1.4. Multi-modal Conversational Emotion Recognition Versus Recurrent Neural Network
The RNN-based emotion recognition methods (Ma et al., 2019; Tao and Liu, 2018; Majumder et al., 2019; Kollias and Zafeiriou, 2020) are developed on the basis of CNN, but they believe that contexts should be mutually influential and interdependent (Ma et al., 2019; Tao and Liu, 2018). These RNN-based methods usually use LSTM or GRU (to avoid gradient disappearance or gradient explosion) to extract semantic features including context. Similar to CNN, RNN can also map emotional features into vector space through mapping functions in an end-to-end manner.
@@ 3.1.5. Multi-modal Conversational Emotion Recognition Versus Transformer
Similar to the RNN-based emotion recognition method, the Transformer-based emotion recognition methods (Huang et al., 2020; Lian et al., 2021; Tsai et al., 2019; Rahman et al., 2020) also extract semantic information including context, and completes subsequent emotion classification based on this (Lian et al., 2021). However, unlike RNN, Transformer’s sequential context modeling ability is better than RNN. Therefore, the accuracy of the Transformer-based emotion recognition methods are significantly better than RNNs.